OpenStudy (anonymous):

What two numbers = 8 when you add them and =63 when you multiply them??

OpenStudy (across):

Let \(x\) and \(y\) be those two numbers. Then\[x+y=8\]and\[xy=63.\]Can you solve this system?

OpenStudy (anonymous):

They have to be the same two numbers.

OpenStudy (across):

That is exactly what this system is trying to do.

OpenStudy (anonymous):

Ok

OpenStudy (across):

\[x+y=8\implies x=8-y\]\[xy=63\implies(8-y)y=63\implies y^2-8y+63=0\implies y=4\pm i\sqrt{47}\]\[x+y=8\implies x+4\pm i\sqrt{47}=8\implies x=4\pm i\sqrt{47}\]